# Sparse Covariance Neural Networks

### v1

@@ 3.2 Soft thresholding
Given the sample covariance matrix $\mathbf{\hat{C}}$ and a coefficient $\tau>0$, we define the soft thresholding function as $\eta(\mathbf{\hat{C}})_{ij}=\hat{c}_{ij}-\textnormal{sign}(\hat{c}_{ij})\tau/\sqrt{t}$ if $|\hat{c}_{ij}|>\tau/\sqrt{t}$, 0 otherwise.
Unlike hard thresholding which only removes small noisy values, soft thresholding subtracts a value from all entries to remove noise also from non-zeroed coefficients. Again, we set the threshold to decrease with the number of samples analogously to (Deshpande & Montanari, 2016) as more accurate covariance estimates reduce the need for sparsification. The following theorem links soft thresholding to the VNN stability.
@@
Consider a soft-thresholded estimate of the covariance matrix $\mathbf{\bar{C}}$ as per Def. 3 with $\tau=M^{\prime}\sqrt{\log(N/c_{0}^{2})}$ and $M^{\prime},C$ two large enough constants. Let the eigenvalues of the true covariance $\{\lambda_{i}\}_{i=0}^{N-1}$ be all distinct. Then, the following holds with high probability:
$\|\mathbf{H}(\mathbf{\bar{C}})\mathbf{x}\!-\!\mathbf{H}(\mathbf{C})\mathbf{x}\|\!\leq\!t^{-1/2}P\sqrt{N}Cc_{0}\max(1,\!\lambda_{\textnormal{max}})\sqrt{\max\left(\log(N/c_{0}^{2}),1\right)}(1+\sqrt{2N})\!+\!\mathcal{O}\left(t^{-1}\!\right)\!.$
Thm. 3 shows that VNNs are stable also when the covariance estimate is soft thresholded and the stability bound decreases with the number of available samples as $t^{-1/2}$. The main takeaways for hard thresholding w.r.t. dense VNNs and sparse PCA hold also for soft-thresholded VNNs. Furthermore, soft thresholding provides better stability than hard thresholding in low-data settings under spiked covariance, as the bound in Thm. 3 contains $\sqrt{\log(N/c_{0}^{2})}$ which is smaller than $\sqrt{\log N}$ in Thm. 2.
Computational complexity. A sparse VNN layer has a computational complexity of order $\mathcal{O}(\|\mathbf{\bar{C}}\|_{0}KF_{\text{in}}F_{\text{out}})$, where $\|\mathbf{\bar{C}}\|_{0}$ is the number of non-zero values of the thresholded sample covariance. This is significantly better than a VNN with dense covariance matrix, which is of order $\mathcal{O}(N^{2}KF_{\text{in}}F_{\text{out}})$, since $\|\mathbf{\bar{C}}\|_{0}$ is generally much smaller than $N^{2}$.
4 Generic True Covariance
Often, the true covariance matrix and its estimate are both dense which makes the VNN computationally heavy. While we can use hard or soft thresholding or any graph sparsification approach, the resulting sparse VNN may be significantly different from the VNN operating on the true covariance, ultimately, leading to divergent outputs and degraded performance. To reduce the computational cost in a tractable manner, we propose a stochastic sparsification framework in a form akin to dropout. Such an approach is general and does not have any structure assumption on the covariance matrix.
@@ 4 Generic True Covariance
Let $\mathbf{\Delta}$ be a matrix of the same support as the sample covariance $\mathbf{\hat{C}}$ with entries $\delta_{ij}=\delta_{ji}=1$ with probability $p_{ij}$ and 0 otherwise (i.e., $\textnormal{Bernoulli}(p_{ij})$) and $\delta_{ii}=1$. A sparsified covariance matrix is $\mathbf{\tilde{C}}=\mathbf{\Delta}\odot\mathbf{\hat{C}}$, where $\odot$ is the elementwise product.
Stochastic sparsification generates random sparsified matrices $\mathbf{\tilde{C}}$ of the sample covariance matrix via element-wise independent sampling in a way that $\mathbf{\tilde{C}}$ preserves: (i) the symmetric property in consistency with the covariance principle; and (ii) the variance of the data points on the main diagonal. The sparsification can be controlled by the probabilities $p_{ij}$ depending on application requirements.
Figure 1: Two-dimensional generalized frequency response $h(\bm{\lambda})$ of stochastic sparsified covariance filter. $h(\bm{\lambda})$ is determined by the coefficients $\{h_{k}\}_{k=0}^{K}$ and is independent of the covariance realizations. For specific covariance realizations, $h(\bm{\lambda})$ is instantiated on the corresponding multivariate frequencies (e.g., $\lambda_{11},\lambda_{12}$ and $\lambda_{21},\lambda_{22}$ for two different realizations $\mathbf{\tilde{C}}_{1}\mathbf{\tilde{C}}_{2}$ and $\mathbf{\tilde{C}}^{\prime}_{1}\mathbf{\tilde{C}}^{\prime}_{2}$).
4.1 VNNs with stochastic covariance sparsification
We now investigate the effects of stochastic sparsification on VNNs’ stability by framing the stochastic sparsification as a stochastic perturbation problem following (Gao et al., 2021a) and defining stochastic sparsified covariance filters.
@@ 4.1 VNNs with stochastic covariance sparsification
Given a sequence of i.i.d. realizations $\mathbf{\tilde{C}}_{k},\dots,\mathbf{\tilde{C}}_{1}$ of the stochastic sparsified covariance $\mathbf{\tilde{C}}$ independent on data distribution, a covariance filter $\mathbf{H}(\mathbf{\tilde{C}})$ performs convolution of a generic signal $\mathbf{x}$ as $\mathbf{\tilde{u}}=\mathbf{H}(\mathbf{\tilde{C}})\mathbf{x}=\sum_{k=0}^{K}h_{k}\mathbf{\tilde{C}}_{k}\dots\mathbf{\tilde{C}}_{1}\mathbf{\tilde{C}}_{0}\mathbf{x}$ with $\mathbf{\tilde{C}}_{0}=\mathbf{I}$.
@@
The generalized frequency response of the stochastic covariance filter is a multivariate function of the form $h(\bm{\lambda})=\sum_{k=0}^{K}h_{k}\prod_{\kappa=0}^{k}\lambda_{\kappa}$ where $\bm{\lambda}=[\lambda_{1},\dots,\lambda_{K}]^{\mathsf{T}}$ with $\lambda_{0}=1$ is a generic vector variable and each frequency variable $\lambda_{k}$ corresponds to the covariance realization $\mathbf{\tilde{C}}_{k}$ for $k=1,\ldots,K$.
The derivations of such a frequency response are reported in Appendix E. Fig. 1 illustrates an example of $K=2$. This extends the deterministic case, in which all realizations and corresponding eigenvalues are equal with $\prod_{\kappa=0}^{k}\lambda_{\kappa}=\lambda^{k}$. Given the generalized frequency response, we generalize the integral Lipschitz property for stability analysis.
@@
Consider the analytic generalized frequency response $h(\bm{\lambda})$ and two instantiations $\bm{\lambda}_{1}=[\lambda_{11},\dots,\lambda_{1K}]^{\mathsf{T}}$ and $\bm{\lambda}_{2}=[\lambda_{21},\dots,\lambda_{2K}]^{\mathsf{T}}$ of vector variable $\bm{\lambda}$. Consider also the auxiliary vector that concatenates the first $k$ entries of $\bm{\lambda}_{2}$ and the last $K-k$ entries of $\bm{\lambda}_{1}$, i.e., $\bm{\lambda}^{(k)}\!=\![\lambda_{21},\dots,\lambda_{2k},\lambda_{1(k+1)},\dots,\lambda_{1K}]^{\mathsf{T}}$. The Lipschitz gradient of $h(\bm{\lambda})$ between $\bm{\lambda}_{1}$ and $\bm{\lambda}_{2}$ is
$\nabla_{L}h(\bm{\lambda}_{1},\bm{\lambda}_{2})=\left[\partial h(\bm{\lambda}^{(1)})/\partial\lambda_{1},\dots,\partial h(\bm{\lambda}^{(K)})/\partial\lambda_{K}\right]^{\mathsf{T}}$
where $\partial h(\bm{\lambda}^{(k)})/\partial\lambda_{k}$ is the partial derivative w.r.t. $\lambda_{k}$ at $\bm{\lambda}^{(k)}$.
@@
A covariance filter is generalized integral Lipschitz if there exists a constant $P$ s.t. $\|\nabla_{L}h(\bm{\lambda}_{1},\bm{\lambda}_{2})\|_{2}\leq P$ and $\|\bm{\lambda}_{1}\odot\nabla_{L}h(\bm{\lambda}_{1},\bm{\lambda}_{2})\|_{2}\leq P.$
The Lipschitz gradient characterizes the variability of $h(\bm{\lambda})$ since, for two multivariate frequency vectors $\bm{\lambda}_{1},\bm{\lambda}_{2}$, we have $h(\bm{\lambda}_{2})-h(\bm{\lambda}_{1})=\nabla_{L}^{\mathsf{T}}h(\bm{\lambda}_{1},\bm{\lambda}_{2})(\bm{\lambda}_{2}-\bm{\lambda}_{1})$. The generalized integral Lipschitz filter limits this variability to be at most linear in the multidimensional space and to decrease as the frequency $\bm{\lambda}$ is specified at large values, which extends the standard integral Lipschitz property to the multivariate frequency domain.
4.2 Stability of VNNs with stochastic covariance sparsification
These preliminaries allow us to analyze the stability of VNNs to stochastic covariance sparsification.
@@ 4.2 Stability of VNNs with stochastic covariance sparsification
Consider a randomly sparsified covariance filter $\mathbf{H}(\mathbf{\tilde{C}})$ [cf. Def. 5] that is generalized integral Lipschitz with constant P [cf. Def. 8]. Let also $\mathbf{H}(\mathbf{C})$ denote the same filter (fixed parameters) operating on the true covariance matrix. Then, for a generic signal $\|\mathbf{x}\|\leq 1$, the expected squared difference between the two filters can be upper-bounded as
@@
$\mathbb{E}[\|\mathbf{H}(\mathbf{C})\mathbf{x}-\mathbf{H}(\mathbf{\tilde{C}})\mathbf{x}\|^{2}]\leq\underbrace{NP^{2}Q+\mathcal{O}((1-p_{1})(1-p_{2}))}_{\textnormal{sparsification error}}+\underbrace{\frac{P^{2}}{t}\mathcal{O}\left(N+\frac{\|\mathbf{C}\|^{2}\log(Nt)}{\nu^{2}t^{2}}\right)}_{\textnormal{covariance uncertainty}}$
@@
where $Q=\sum_{i=1}^{N}\sum_{n=1}^{N}\hat{c}_{in}^{2}(1-p_{in})$, $p_{1},p_{2}$ are generic probabilities such that $\mathcal{O}((1-p_{1})(1-p_{2}))$ is dominated by the linear terms in the probability value in $Q$ and $\nu$ is defined in Thm. 1.
@@
Thm. 4 identifies two main factors that affect the stability of stochastic sparsified covariance filters: the covariance uncertainty and the sparsification error.
Covariance uncertainty. This term is analogous to the dense case in Thm. 1. It decreases with the number of samples as $\mathcal{O}(1/t)$ and is dominated by the sparsification error for sufficiently large $t$. Note that a large $t$ does not necessarily correspond to a small covariance estimation error here because the stochastic sparsification may still remove large covariance values; ultimately, making our analysis principally different from (Sihag et al., 2022) which assumes small covariance perturbations.
@@
Sparsification error. This term decreases as the sampling probabilities $p_{ij}\to 1$, corresponding to an improved stability but a lower sparsification. This indicates a trade-off between the perturbation effect caused by the covariance sparsification error and the computational cost saved by the stochastic sparsification. The stability constant depends on the filter Lipschitz constant $P$, the data dimension $N$, and the coupling between the covariance values $\hat{c}_{ij}$ and the sampling probabilities $p_{ij}$, i.e., $Q$. A larger $P$ allows for a higher filter discriminability as the frequency response can change more quickly but leads to worse stability. Data with high dimensionality $N$ results in a larger graph and increases the effect of sparsification on the VNN stability. More importantly, $Q$ represents the interplay between sample covariance values $\hat{c}_{ij}$ and their corresponding sampling probabilities $p_{ij}$, which can be used as a design choice to develop different stochastic sparsification strategies in a principled manner as we shall discuss in the next section. Finally, note that the results in Thm. 4 and its proof generalize and differ substantially from the results in (Gao et al., 2021a, Theorem 1) as we consider here edge-specific probabilities $p_{ij}$ rather than an identical probability $p_{ij}=p$ and we identify a connection between VNN stability and the data distribution through their covariances $\hat{c}_{ij}$, which is not present in (Gao et al., 2021a).
4.3 Stochastic sparsification strategies
Following our theoretical analysis, we propose two strategies to assign the sparsification probabilities to covariance values and assess their impact on the stability of VNN.

### v2

@@ III-B S-VNNs with Soft Thresholding
Given the sample covariance matrix $\mathbf{\hat{C}}$ and a coefficient $\tau>0$, we define the soft thresholding function as $\eta(\mathbf{\hat{C}})_{ij}=\hat{c}_{ij}-\textnormal{sign}(\hat{c}_{ij})\tau/\sqrt{t}$ if $|\hat{c}_{ij}|>\tau/\sqrt{t}$, and 0 otherwise.
Unlike hard thresholding, which only removes small uncertain values, soft thresholding subtracts a value from all entries to remove uncertainty also from the non-zeroed coefficients. Again, we set the threshold to decrease with the number of samples $t$ analogously to [18] as more accurate covariance estimates reduce the need for sparsification.
@@
Consider a soft-thresholded estimate of the covariance matrix $\mathbf{\bar{C}}$ as per Def.˜3 with $\tau=M^{\prime}\sqrt{\log(N/c_{0}^{2})}$ and $M^{\prime},C$ two large enough constants and consider a covariance filter $\mathbf{H}(\cdot)$ that is Lipschitz with constant $P$. Let the eigenvalues of the true covariance $\{\lambda_{i}\}_{i=1}^{N}$ be all distinct and strictly positive. Then, the following holds with probability 1o(1)1-o(1):
$\|\mathbf{H}(\mathbf{\bar{C}})\mathbf{x}\!-\!\mathbf{H}(\mathbf{C})\mathbf{x}\|\!\leq\!\mathcal{O}\left(t^{-1}\!\right)\!+$
$t^{-\frac{1}{2}}P\sqrt{N}Cc_{0}\max(1,\!\lambda_{\textnormal{max}})\sqrt{\max\left(\log(N/c_{0}^{2}),1\right)}(1+\sqrt{N}).$
Thm.˜4 shows that S-VNNs are stable also when the covariance estimate is soft thresholded and the stability bound decreases with the number of samples at the rate of $t^{-1/2}$. The main takeaways for hard thresholded S-VNNs w.r.t. VNNs and sparse PCA hold also for soft-thresholded S-VNNs. Moreover, soft thresholding achieves a better stability than hard thresholding in low-data settings as the term $\sqrt{\log(N/c_{0}^{2})}$ in Thm. 4 is smaller than the term $\sqrt{\log N}$ in Thm. 3.
Computational complexity. An S-VNN layer has a computational complexity of order $\mathcal{O}(\|\mathbf{\bar{C}}\|_{0}KF_{\text{in}}F_{\text{out}})$, where $\|\mathbf{\bar{C}}\|_{0}$ is the number of non-zero values of the thresholded sample covariance. This is significantly better than a VNN with a computational complexity of order $\mathcal{O}(N^{2}KF_{\text{in}}F_{\text{out}})$, since $\|\mathbf{\bar{C}}\|_{0}$ is generally much smaller than $N^{2}$.
IV Generic True Covariance
In this section, we consider the more general and challenging setup where the true covariance matrix and consequently its estimate are dense; which makes the computation of VNNs heavy. While thresholding strategies can be applied to this setting, choosing an appropriate threshold is challenging; a too large threshold might lead to an S-VNN that is significantly different from the VNN operating on the true covariance due to largely truncated correlations, ultimately leading to decreases in performance, while a too small threshold might lead to an irrelevant computational advantage, thus defeating the purpose of sparsification. To overcome this issue and reduce computational cost in a tractable manner, we propose a stochastic sparsification framework in a form akin to dropout. Such an approach is general and does not require any structure assumption on the true covariance matrix, and it also extends the former stability analysis from small to large perturbations as stochastic sparsification may lead to large changes on the sample covariance matrix.
@@ IV Generic True Covariance
Let $\mathbf{\Delta}$ be a matrix with the same support as the sample covariance $\mathbf{\hat{C}}$, with entries $\delta_{ij}=\delta_{ji}=1$ with probability $p_{ij}$ and 0 otherwise (i.e., $\textnormal{Bernoulli}(p_{ij})$) and $\delta_{ii}=1$. A sparsified covariance matrix is $\mathbf{\tilde{C}}=\mathbf{\Delta}\odot\mathbf{\hat{C}}$, where $\odot$ is the elementwise product.
Stochastic sparsification generates randomly sparsified matrices of the sample covariance matrix via element-wise independent sampling, and allows the sparsified matrix $\mathbf{\tilde{C}}$ to preserve: (i) the symmetric property in consistency with the covariance principle; and (ii) the variance of the data points on the main diagonal.
Figure 3: Two-dimensional generalized frequency response $h(\boldsymbol{\lambda})$ of stochastic sparsified covariance filter. $h(\boldsymbol{\lambda})$ is determined by the coefficients $\{h_{k}\}_{k=0}^{K}$ and is independent of the covariance realizations. For specific realizations, $h(\boldsymbol{\lambda})$ is instantiated on the corresponding multivariate frequencies (e.g., $\lambda_{11},\lambda_{12}$ and $\lambda_{21},\lambda_{22}$ for two different realizations $\mathbf{\tilde{C}}_{1}\mathbf{\tilde{C}}_{2}$ and $\mathbf{\tilde{C}}^{\prime}_{1}\mathbf{\tilde{C}}^{\prime}_{2}$).
IV-A S-VNNs with Stochastic Sparsification
We now investigate the effects of stochastic sparsification on the S-VNN by formulating a stochastic perturbation problem following [36]. We start by defining stochastic covariance filters.
@@ IV-A S-VNNs with Stochastic Sparsification
[…]
@@
The generalized frequency response of the stochastic covariance filter is a multivariate function of the form $h(\boldsymbol{\lambda})=\sum_{k=0}^{K}h_{k}\prod_{\kappa=0}^{k}\lambda_{\kappa}$ where $\lambda_{0}=1$ and $\boldsymbol{\lambda}=[\lambda_{1},\dots,\lambda_{K}]^{\mathsf{T}}$ is a generic vector variable with each frequency variable $\lambda_{k}$ corresponding to the covariance realization $\mathbf{\tilde{C}}_{k}$ for $k=1,\ldots,K$.
The derivations of such a frequency response are reported in Appendix D in the supplement. Intuitively, the generalized frequency response is a multivariate function in the eigenvalues of each random covariance realization $\mathbf{\tilde{C}}_{k}$ at each shift $k$, whereas for a fixed $\mathbf{\hat{C}}$, the eigenvalues are the same for each shift and therefore the resulting function is a univariate polynomial. Fig.˜3 illustrates an example of $K=2$. This extends the covariance filter frequency response in Sec. II-B to the stochastic setting.
Given the generalized covariance filter frequency response, we generalize the integral Lipschitz property for stability analysis.
@@
Consider the analytic generalized frequency response $h(\boldsymbol{\lambda})$ and two instantiations $\boldsymbol{\lambda}_{1}=[\lambda_{11},\dots,\lambda_{1K}]^{\mathsf{T}}$ and $\boldsymbol{\lambda}_{2}=[\lambda_{21},\dots,\lambda_{2K}]^{\mathsf{T}}$ of vector variable $\boldsymbol{\lambda}$. Consider also the auxiliary vector that concatenates the first $k$ entries of $\boldsymbol{\lambda}_{2}$ and the last KkK-k entries of $\boldsymbol{\lambda}_{1}$, i.e., $\boldsymbol{\lambda}^{(k)}\!=\![\lambda_{21},\dots,\lambda_{2k},\lambda_{1(k+1)},\dots,\lambda_{1K}]^{\mathsf{T}}$. The Lipschitz gradient of $h(\boldsymbol{\lambda})$ between $\boldsymbol{\lambda}_{1}$ and $\boldsymbol{\lambda}_{2}$ is
$\nabla_{L}h(\boldsymbol{\lambda}_{1},\boldsymbol{\lambda}_{2})=\left[\partial h(\boldsymbol{\lambda}^{(1)})/\partial\lambda_{1},\dots,\partial h(\boldsymbol{\lambda}^{(K)})/\partial\lambda_{K}\right]^{\mathsf{T}}$
where $\partial h(\boldsymbol{\lambda}^{(k)})/\partial\lambda_{k}$ is the partial derivative w.r.t. $\lambda_{k}$ at $\boldsymbol{\lambda}^{(k)}$.
@@
A covariance filter is generalized integral Lipschitz if there exists a constant $P$ s.t. $\|\nabla_{L}h(\boldsymbol{\lambda}_{1},\boldsymbol{\lambda}_{2})\|\leq P$ and $\|\boldsymbol{\lambda}_{1}\odot\nabla_{L}h(\boldsymbol{\lambda}_{1},\boldsymbol{\lambda}_{2})\|\leq P.$
The Lipschitz gradient characterizes the variability of $h(\boldsymbol{\lambda})$ since, for two multivariate frequency vectors $\boldsymbol{\lambda}_{1},\boldsymbol{\lambda}_{2}$, we have $h(\boldsymbol{\lambda}_{2})-h(\boldsymbol{\lambda}_{1})=\nabla_{L}^{\mathsf{T}}h(\boldsymbol{\lambda}_{1},\boldsymbol{\lambda}_{2})(\boldsymbol{\lambda}_{2}-\boldsymbol{\lambda}_{1})$. The generalized integral Lipschitz filter limits this variability to be at most linear in the multidimensional space and to decrease as the frequency $\boldsymbol{\lambda}$ is specified at large values, which extends the standard integral Lipschitz property to the multivariate frequency domain.
These preliminaries allow us to analyze the stability of S-VNNs with stochastic covariance sparsification.
@@
[…]
@@
$\mathbb{E}[\|\mathbf{H}(\mathbf{C})\mathbf{x}-\mathbf{H}(\mathbf{\tilde{C}})\mathbf{x}\|^{2}]\leq\underbrace{NP^{2}Q+\mathcal{O}((1-p_{1})(1-p_{2}))}_{\textnormal{sparsification error}}+$
$\underbrace{\frac{P^{2}k_{\textnormal{max}}^{2}}{t^{1-2\epsilon}}\mathcal{O}\left(N+\frac{\|\mathbf{C}\|^{2}\log(Nt)}{k_{\textnormal{min}}^{2}t^{4\epsilon}}\right)}_{\textnormal{covariance uncertainty}}$
@@
where $Q=\sum_{i=1}^{N}\sum_{n=1}^{N}\hat{c}_{in}^{2}(1-p_{in})$, the residual $\mathcal{O}((1-p_{1})(1-p_{2}))$ for generic probababilities $p_{1},p_{2}$ collects higher-order probability terms that are negligible compared the linear contributions in the probability value in $Q$, and $k_{\textnormal{max}},k_{\textnormal{min}},\kappa$ are defined in Thm.˜1.
@@
Thm.˜5 identifies two main factors that affect the stability of sparse covariance filters with stochastic sparsification, i.e., the covariance uncertainty and the sparsification error.
Covariance uncertainty. This term is analogous to the stability bound of VNN in Thm.˜1. It decreases with the number of samples as $\mathcal{O}(1/{t^{1-2\epsilon}})$ such that the overall bound is dominated by the sparsification error for sufficiently large $t$ (i.e., the error introduced by sparsification prevails on the finite-sample estimation error).
@@
Sparsification error. This term decreases as the sampling probabilities $p_{ij}\to 1$, corresponding to an improved stability but a lower sparsification. This indicates a trade-off between the perturbation effect caused by the covariance sparsification error and the computational cost saved by the stochastic sparsification. The stability constant depends on the filter Lipschitz constant $P$, the data dimension $N$, and the coupling between the covariance values $\hat{c}_{ij}$ and the sampling probabilities $p_{ij}$, i.e., $Q$. A larger $P$ allows for a higher filter discriminability as the frequency response can change more quickly and identify the difference between nearby frequencies, but leads to worse stability. Data with high dimensionality $N$ results in a larger graph and increases the effect of sparsification on the VNN stability. More importantly, $Q$ represents the interplay between sample covariance values $\hat{c}_{ij}$ and their corresponding sampling probabilities $p_{ij}$, which can be used as a design choice to develop stochastic sparsification strategies in a principled manner as we shall discuss next. This term allows to capture the impact of large perturbations caused by sparsification, making our analysis principally different from [11] which is limited to small covariance perturbations. Finally, note that the results in Thm. 5 and its proof generalize and differ substantially from the results in [36, Theorem 1] as we consider here edge-specific probabilities $p_{ij}$ rather than an identical probability $p_{ij}=p$ and we identify a connection between the S-VNN stability and the data distribution through their covariances $\hat{c}_{ij}$, which is not present in [36].
Figure 4: Insights into stability for stochastic sparsification. (Left) Stability bound elements $q_{in}=\hat{c}_{in}^{2}(1-|\hat{c}_{in}|/\hat{c}_{\text{max}})$ with $Q=\sum_{i=1}^{N}\sum_{n=1}^{N}q_{in}$ for ACV [cf. Thm.˜5]. (Right) Stability term $Q$ [cf. Thm.˜5] for covariance values distributions with different means and stochastic sparsifications.
IV-B Sparsification Probability
Following the interplay between the sample covariance values and sparsification probabilities on the S-VNN stability, we propose two stochastic sparsification strategies termed: absolute covariance values and ranked covariance values.